OpenStudy (anonymous):

What is the value of x-y in this expression ? 2(x+2/5) = 2 (3/7) +2y ?

OpenStudy (anonymous):

2x + 4/5 = 6/7 +2y 2x - 2y = 6/7 - 4/5 x-y = 3/7 - 2/5 =1/35

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You are welcome.